# Algorithmic stability implies training-conditional coverage for distribution-free prediction methods

### v2

@@ Appendix A Additional proofs
A.3 Proof of Theorem 4.2 (full conformal—alternative version)
@@
A.4 Proof of Theorem 4.5 (full conformal with swap-stability)
@@
For the finite-sample guarantee, on the other hand, the proof for jackknife+, which relied on the “weak converse” relating out-of-sample stability to out-of-sample swap-stability (Proposition 4.6), cannot be reproduced for full conformal, because the “weak converse” result may no longer hold for in-sample stability; we will need to take a different approach.
@@
A.5 Proof of Proposition 5.1
@@
Now define $A=\big{(}(X_{1},Y_{1}),\dots,(X_{n-m^{\prime}},Y_{n-m^{\prime}})\big{)}$, $B=\big{(}(X_{n-m^{\prime}+1},Y_{n-m^{\prime}+1}),\dots,(X_{n-m},Y_{n-m})\big{)}$, and $C=\big{(}(X_{n-m+1},Y_{n-m+1}),\dots,(X_{n},Y_{n})\big{)}$. Then, following the notation of Lemma A.1, we can write
@@
for the appropriately defined function $f$ (i.e., $f$ is the function that applies the algorithm $\mathcal{A}$ to all $n$ data pairs contained in $A,B,C$, and then evaluates the resulting fitted model on the last component of $A$). Applying the lemma immediately verifies that $\bar{\beta}^{\textnormal{out}}_{m,n}(\mathcal{A},P)\leq\bar{\beta}^{\textnormal{out}}_{m^{\prime},n}(\mathcal{A},P)$.
@@
Appendix B Proofs for $m$-stability calculations
B.1 Proof of Proposition 5.3 ($k$-nearest neighbors)
@@ Appendix B Proofs for $m$-stability calculations
B.2 Proof of Proposition 5.4 (ridge regression)
@@
$\mathbb{E}\left[\textnormal{(Term 1)}\right]\leq\frac{\lambda^{-1}m}{n+m}\mathbb{E}\left[\|\widehat{S}_{>n}-\widehat{S}_{n}\|_{2}\right]\\ \leq\frac{\lambda^{-1}m}{n+m}\left(\mathbb{E}\left[\|\widehat{S}_{>n}-\mathbb{E}[XY]\|_{2}\right]+\mathbb{E}\left[\|\widehat{S}_{n}-\mathbb{E}[XY]\|_{2}\right]\right)\\ \leq\frac{\lambda^{-1}m}{n+m}\left(\mathbb{E}\left[\|\widehat{S}_{>n}-\mathbb{E}[XY]\|^{2}_{2}\right]^{1/2}+\mathbb{E}\left[\|\widehat{S}_{n}-\mathbb{E}[XY]\|^{2}_{2}\right]^{1/2}\right).$
@@
$\left(\widehat{\Sigma}_{n}+\frac{m}{n+m}(\widehat{\Sigma}_{>n}-\widehat{\Sigma}_{n})+\lambda\mathbf{I}_{d}\right)^{-1}-\left(\widehat{\Sigma}_{n}+\lambda\mathbf{I}_{d}\right)^{-1}\\ =-\left(\widetilde{\Sigma}_{t}+\lambda\mathbf{I}_{d}\right)^{-1}\cdot\left[\frac{m}{n+m}(\widehat{\Sigma}_{>n}-\widehat{\Sigma}_{n})\right]\cdot\left(\widetilde{\Sigma}_{t}+\lambda\mathbf{I}_{d}\right)^{-1}$
@@
B.3 Proof of Proposition 5.5 (bagging)
@@
for any $n$. (We can view this as a stronger version of the out-of-sample $1$-stability condition—that is, this bounds the expected squared perturbation, i.e., the $L_{2}$ norm rather than the $L_{1}$ norm.) Replacing $n$ with $n+k-1$ (for some $1\leq k\leq m$), then,
@@
(Here, and below, for each $k\geq 0$ we define the model $\widetilde{\mu}_{n+k}=\widetilde{\mathcal{A}}_{\textnormal{bag}}(\mathcal{D}_{n+k})$, i.e., the output of derandomized bagging run on the data set $\mathcal{D}_{n+k}=\big{\{}(X_{i},Y_{i})\big{\}}_{i\in[n+k]}$, where $(X_{1},Y_{1}),(X_{2},Y_{2}),\dots$ denotes a sequence of data points sampled i.i.d. from $P$.)
@@
from the calculations above. Next we will verify that the second term is zero. Assume $k<k^{\prime}$ without loss of generality. We need to verify that $\mathbb{E}\left[\left(\widetilde{\mu}_{n+k-1}(X)-\widetilde{\mu}_{n+k}(X)\right)\cdot\left(\widetilde{\mu}_{n+k^{\prime}-1}(X)-\widetilde{\mu}_{n+k^{\prime}}(X)\right)\right]=0$. First, for any $i\in[n+k]$, define
@@
In other words, the average $\frac{1}{n+k}\sum_{i=1}^{n+k}\widetilde{\mu}_{[n+k]\backslash i}(X)$ is constructed by randomly sampling $i\in[n+k]$, and then randomly sampling a subset $S\subseteq[n+k]\backslash i$; this is equivalent to simply sampling $S\subseteq[n+k]$, i.e., computing $\widetilde{\mu}_{n+k}(X)$. Therefore, $\left(\frac{1}{n+k}\sum_{i=1}^{n+k}\widetilde{\mu}_{[n+k]\backslash i}(X)-\widetilde{\mu}_{n+k}(X)\right)=0$ almost surely. This verifies that $\mathbb{E}\left[\left(\widetilde{\mu}_{n+k-1}(X)-\widetilde{\mu}_{n+k}(X)\right)\cdot\left(\widetilde{\mu}_{n+k^{\prime}-1}(X)-\widetilde{\mu}_{n+k^{\prime}}(X)\right)\right]=0$ for any $k\neq k^{\prime}$.

### v3

@@ Appendix B Proofs of training-conditional coverage under alternative stability frameworks
B.2 Proof of Theorem 4.2 (full conformal—alternative version)
@@
B.3 Proof of Theorem 4.4 (jackknife+ with swap-stability)
As mentioned in Section 4.2, the swap-stability is sufficient for jackknife+ to achieve training-conditional coverage because a weaker version of converse to Proposition 4.3 can be established for out-of-sample stability. To state the partial converse result, we introduce some notations. Recalling that an algorithm $\mathcal{A}$ maps a data set of any size to a fitted function, we will write $\mathcal{A}_{n}$ to denote the map defined by applying $\mathcal{A}$ to a training set of size $n$. That is, $\mathcal{A}_{n}$ is a map with input space $(\mathcal{X}\times\mathbb{R})^{n}$, and $\mathcal{A}$ is simply the collection of maps $\mathcal{A}_{n}$ across all $n$.
Proposition B.1.
Proof of Proposition B.1.
Now we have
Proof of Theorem 4.4 (jackknife+ with swap-stability).
First, we can observe that $\widehat{C}_{n}^{\textnormal{$\gamma$-J+}}(X_{n+1})$ depends on the algorithm $\mathcal{A}$ only through the map $\mathcal{A}_{n-1}$; that is, when running jackknife+, $\mathcal{A}$ is only ever trained on data sets of size $n-1$. Now let $\mathcal{A}^{\prime}$ be defined as in Proposition B.1 (except with $n-1$ in place of $n$), so that we have $\mathcal{A}^{\prime}_{n-1}=\mathcal{A}_{n-1}$. In other words, at sample size $n-1$, running jackknife+ with $\mathcal{A}$ is exactly equivalent to running jackknife+ with algorithm $\mathcal{A}^{\prime}$. Therefore, the results of Theorems 3.2 and 3.3 can be applied with $\mathcal{A}^{\prime}$ in place of $\mathcal{A}$; applying Proposition B.1 tells us that $\beta_{m,n-1}^{\textnormal{out}}(\mathcal{A}^{\prime},P)\leq\bar{\beta}_{m,n-1}^{\textnormal{out}}(\mathcal{A},P)$, which verifies the desired finite-sample training-coverage guarantee. Finally, the asymptotic guarantees with swap-stability are proved exactly as for (non-swap) stability, i.e., following the same arguments as in the proof of Theorem 3.6. ∎
B.4 Proof of Theorem 4.5 (full conformal with swap-stability)
@@
For the finite-sample guarantee, on the other hand, the proof for jackknife+, which relied on the “weak converse” relating out-of-sample stability to out-of-sample swap-stability (Proposition B.1), cannot be reproduced for full conformal, because the “weak converse” result may no longer hold for in-sample stability; we will need to take a different approach.
@@
Appendix C Proofs for $m$-stability calculations
C.1 Proof of Proposition 5.1 (monotonicity of swap-stability in $m$)
@@ Appendix C Proofs for $m$-stability calculations
Now define
$A$ $=\big{(}(X_{1},Y_{1}),\dots,(X_{n-m^{\prime}},Y_{n-m^{\prime}})\big{)},$
$B$ $=\big{(}(X_{n-m^{\prime}+1},Y_{n-m^{\prime}+1}),\dots,(X_{n-m},Y_{n-m})\big{)},$
$C$ $=\big{(}(X_{n-m+1},Y_{n-m+1}),\dots,(X_{n},Y_{n})\big{)}.$
Then, following the notation of Lemma C.1, we can write
@@
for the appropriately defined function $f$ (i.e., $f$ is the function that applies the algorithm $\mathcal{A}$ to all $n$ data pairs contained in $A,B,C$, and then evaluates the resulting fitted model on an independent copy of the data $(X,Y)$). Applying the lemma immediately verifies that $\bar{\beta}^{\textnormal{out}}_{m,n}(\mathcal{A},P)\leq\bar{\beta}^{\textnormal{out}}_{m^{\prime},n}(\mathcal{A},P)$.
@@
C.2 Proof of Lemma 5.2 (relating $1$-stability and $m$-stability)
by definition of out-of-sample $1$-swap-stability together with the assumption of symmetry of $\mathcal{A}$. Again, a similar argument holds for in-sample swap-stability.
C.3 Proof of Proposition 5.3 ($k$-nearest neighbors)
@@
C.4 Proof of Proposition 5.4 (ridge regression)
@@
$\mathbb{E}\left[\textnormal{(Term 1)}\right]\leq\frac{\lambda^{-1}m}{n+m}\mathbb{E}\left[\|\widehat{S}_{>n}-\widehat{S}_{n}\|_{2}\right]\leq\frac{\lambda^{-1}m}{n+m}\left(\mathbb{E}\left[\|\widehat{S}_{>n}-\mathbb{E}[XY]\|_{2}\right]+\mathbb{E}\left[\|\widehat{S}_{n}-\mathbb{E}[XY]\|_{2}\right]\right)\\ \leq\frac{\lambda^{-1}m}{n+m}\left(\mathbb{E}\left[\|\widehat{S}_{>n}-\mathbb{E}[XY]\|^{2}_{2}\right]^{1/2}+\mathbb{E}\left[\|\widehat{S}_{n}-\mathbb{E}[XY]\|^{2}_{2}\right]^{1/2}\right).$
@@
$\left(\widehat{\Sigma}_{n}+\frac{m}{n+m}(\widehat{\Sigma}_{>n}-\widehat{\Sigma}_{n})+\lambda\mathbf{I}_{d}\right)^{-1}-\left(\widehat{\Sigma}_{n}+\lambda\mathbf{I}_{d}\right)^{-1}=-\left(\widetilde{\Sigma}_{t}+\lambda\mathbf{I}_{d}\right)^{-1}\cdot\left[\frac{m}{n+m}(\widehat{\Sigma}_{>n}-\widehat{\Sigma}_{n})\right]\cdot\left(\widetilde{\Sigma}_{t}+\lambda\mathbf{I}_{d}\right)^{-1}$
@@
C.5 Proof of Proposition 5.5 (bagging)
@@
for any $n$. We can view this as a stronger version of the out-of-sample $1$-stability condition—that is, this bounds the expected squared perturbation, i.e., the $L_{2}$ norm rather than the $L_{1}$ norm. Replacing $n$ with $n+k-1$ (for some $1\leq k\leq m$), then,
@@
Here, and below, for each $k\geq 0$ we define the model $\widetilde{\mu}_{n+k}=\widetilde{\mathcal{A}}_{\textnormal{bag}}(\mathcal{D}_{n+k})$, i.e., the output of derandomized bagging run on the data set $\mathcal{D}_{n+k}=\big{\{}(X_{i},Y_{i})\big{\}}_{i\in[n+k]}$, where $(X_{1},Y_{1}),(X_{2},Y_{2}),\dots$ denotes a sequence of data points sampled i.i.d. from $P$.
@@
from the calculations above. Next we will verify that the second term is zero. Assume $k<k^{\prime}$ without loss of generality. We need to verify that
$\mathbb{E}\left[\left(\widetilde{\mu}_{n+k-1}(X)-\widetilde{\mu}_{n+k}(X)\right)\cdot\left(\widetilde{\mu}_{n+k^{\prime}-1}(X)-\widetilde{\mu}_{n+k^{\prime}}(X)\right)\right]=0.$
First, for any $i\in[n+k]$, define
@@
In other words, the average $\frac{1}{n+k}\sum_{i=1}^{n+k}\widetilde{\mu}_{[n+k]\backslash i}(X)$ is constructed by randomly sampling $i\in[n+k]$, and then randomly sampling a subset $S\subseteq[n+k]\backslash i$; this is equivalent to simply sampling $S\subseteq[n+k]$, i.e., computing $\widetilde{\mu}_{n+k}(X)$. Therefore,
$\left(\frac{1}{n+k}\sum_{i=1}^{n+k}\widetilde{\mu}_{[n+k]\backslash i}(X)-\widetilde{\mu}_{n+k}(X)\right)=0$
almost surely. This verifies that $\mathbb{E}\left[\left(\widetilde{\mu}_{n+k-1}(X)-\widetilde{\mu}_{n+k}(X)\right)\cdot\left(\widetilde{\mu}_{n+k^{\prime}-1}(X)-\widetilde{\mu}_{n+k^{\prime}}(X)\right)\right]=0$ for any $k\neq k^{\prime}$.